OpenStudy (anonymous):

Complex Fractions.... How do I simplify 2 - 8 over X+4 divided by 3 - 12 over X+4?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

The answer is two thirds but I confused with the steps.

OpenStudy (anonymous):

me to

OpenStudy (anonymous):

[(2 - 8) / (X+4)] / [(3 - 12) / (X+4)] = [-6 / (X+4)] / [-9 / (X+4)] = [-6 / (X+4)] * [(X+4) / -9] The (x+4)s cancal out, so you have -6/-9 = 2/3

OpenStudy (anonymous):

just look in google

OpenStudy (anonymous):

the 4x cancels out the X+4?

OpenStudy (anonymous):

no x-4 cancels out with x-4

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

I have another question, what is 8/x+4 times 8x?

OpenStudy (anonymous):

[8/(x+4)] * 8x or (8/x) + 4 *8x ?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

[8/(x+4)] * 8x = 64x / (x+4)[8/(x+4)] * 8x = 64x / (x+4) You can't really simplify further

OpenStudy (anonymous):

whoops ignore the second line

OpenStudy (anonymous):

ahh still confused methods haha thanks a bunch!

OpenStudy (anonymous):

no problem, good luck!